OpenStudy (anonymous):

5^x=3^x+1 how to solve that. I know i have to get a common base, but i don't know how?

OpenStudy (anonymous):

you want to use logarithms. because of the laws of logarithms the exponent can be brought down.

OpenStudy (anonymous):

\[\large \log a^b=b \log a\]

OpenStudy (anonymous):

take the log of both sides would be step 1

OpenStudy (anonymous):

would it be xlog base 5 = x+1log base 3

OpenStudy (anonymous):

\[x \ln(5)=x \ln(3) +1\]\[x (\ln(5) - x \ln(3) =1\]\[x ( \ln(5)-\ln (1) = 1\]\[x= \frac{1}{\ln(5) - \ln ( 3)}\]use calculator solve ,the rest is your

OpenStudy (anonymous):

log has a base of 10 unless specified otherwise. \[\log_{10} 1=x \rightarrow 10^x=1\]

OpenStudy (anonymous):

sorry ln (3) not ln(1)

OpenStudy (anonymous):

third row I was wrong

OpenStudy (beginnersmind):

ChmE log(3^x +1) isn't equal to log(3^x) + log 1

OpenStudy (anonymous):

log 1 is 0 because anything raised to the zero. ln 1 is also equal to zero.\[\ln_e 1 = x \rightarrow e^x=1 \rightarrow x=0\]

zepdrix (zepdrix):

Was the +1 in the exponent or no?

zepdrix (zepdrix):

\[\huge 5^x=3^{x+1}\]Like this?

OpenStudy (anonymous):

yea

zepdrix (zepdrix):

Then yes you had the right idea with your response before. But you have to take the SAME LOG of both sides, let's take log base 5 of both sides,\[\huge \log_5 5^x=\log_5 3^{x+1}\]And now you can do that little tricky you mentioned before,\[\huge x \log_5 5=(x+1) \log_5 3\]

zepdrix (zepdrix):

This is a good problem, it's a little tougher than the rest you've been doing. If you can get through this one, then you've got a good handle on the material.

OpenStudy (anonymous):

i was told to get a common base, so both sides have to be the same and then they cancel out and the x and the x+1 comes down and you solve for the answer

zepdrix (zepdrix):

Hmm :o

OpenStudy (anonymous):

you should write equation in racket

OpenStudy (anonymous):

thats why i couldn't get it, since i couldn't get the bases to be the same

OpenStudy (beginnersmind):

You can write the LHS as \[(3^{log_{3}5})^{x}\]

OpenStudy (watchmath):

\(5^x=3^{x+1}=3\cdot 3^x\) \((5/3)^x=3\) \(x\ln(5/3)=\ln 3\) \(x=\frac{\ln 3}{\ln(5/3)}\)

zepdrix (zepdrix):

Ok so perhaps you're learning it using a different method in your class. I'll just explain this method in case it helps.\[\large 5^x=3^{x+1}\]Taking the natural log of both sides,\[\large \ln 5^x = \ln 3^{x+1}\]Which can be written as,\[\large x \ln5=(x+1) \ln 3\]Distribute the ln3 to each term in the brackets,\[\large x \ln5=x \ln3+ \ln3\]Subtract x ln3 from each side,\[\large x \ln5-x \ln3=\ln 3\]Factoring out an x,\[\large x(\ln5-\ln3)=\ln 3\] \[\large x\quad =\quad \frac{\ln3}{\ln5-\ln3}\quad =\quad \frac{\ln 3}{\ln(\frac{5}{3})}\] It looks like watch's method is probably more of what you were looking for though.

OpenStudy (anonymous):

oh ok. i get it now. thank you.